OpenStudy (anonymous):

what number is next? -7, 0, 19, 56

OpenStudy (anonymous):

117

OpenStudy (anonymous):

am i right???

OpenStudy (anonymous):

Care to justify your claim? I see no pattern.

OpenStudy (anonymous):

i believe so...how did you get your answer?

OpenStudy (anonymous):

hm does this have something to do with some kinda primes or something?

OpenStudy (anonymous):

first two have a difference of 7 2nd and 3rd have 19 then 37 Now take secondary series as 7,19,37 differnce is 12,18 so next difference would be 24 hence next no. in second series is 24+37=61 and hence next number in first series is 56+61=117

OpenStudy (anonymous):

Seems legit.

OpenStudy (asnaseer):

sahil is correct, you could also spot the sequence as \(n^3-8\) n=1,2,3,...

OpenStudy (anonymous):

oh I couldn't figure that out

OpenStudy (anonymous):

how about -1, 6, 25, 62

OpenStudy (anonymous):

7 19 37 are the differences here... same as last time. Refer to the last problem to figure out what to do next.

OpenStudy (asnaseer):

mrs237 - I see you are new to OpenStudy since this is your first question - Welcome aboard! The normal protocol here is to ask one question, wait for answers and give a medal to the best answer (in your opinion) if you want to. Then close the question and ask another. So please post your new question by first closing this one. Thank you.